HUBBLE SPACE TELESCOPE DAILY REPORT #925 PERIOD COVERED: 0800 EST Daily Status Report as of 190/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 5016 (The Evolution of the UV Spectra in Early Type Galaxies Out to Z = 0.7) The Wide Field Camera was used to make four images of the weak radio galaxy 53W034 to study its current morphology. This study will be later followed by proposal 3545 to do a complete spectrographic study of the galaxy. The first three observations were received in real time, and were reported to look okay. There was a recentering event which occurred during the third observation. The fourth observation occurred during LOS and will be analyzed later. 1.2 Completed FOS/RD 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle 3 High) The Faint Object Spectrograph was used to make four observations of quasar 1148+5454 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observations were received and displayed. The observations were reported to be successful, and to look fine. 1.3 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make a snapshot image of the protoplanetary system WX-PSC to study its current morphology. One observations was received and examined. No anomalies were noted. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.5 Completed FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) The Faint Object Spectrograph was used to make observations of quasar 2349-0125 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The target acquisition was successful, and the observations was reported to look good. 1.6 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats, Narrow 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled, and no problems were reported. 1.7 Completed FOC/96 4113 (Auroral Imaging of Jupiter With the FOC Cycle 2, February- March 1993) The Faint Object Camera was used to make two images of the auroral emissions on the planet Jupiter at 1600 A0. Both north and south observations were contained within the sequence. After verifying that the correct filter were in place, the OSS successfully uplinked two corrective SIP maneuvers. No problems were reported with the images. 1.8 Completed FOS/BL 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of quasar 0448-3916 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.9 Completed FOS/BL 3618 (Excitation Processes for the Outer Planet UV Emissions: Future - Cycle C) The Faint Object Spectrograph (Blue Side) was used to observe the planet Saturn to learn more about the processes responsible for the "electroglow" emission in the planet's upper atmosphere. Blind pointing was used to acquire the target and course track was used for guiding, since diffuse emission was being observed and the proposes want to avoid loss of lock. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Note: The two day long Rendezvous and Grapple Simulation was completed yesterday (i.e., covering Wednesday and Thursday, July 7-8). All major objective of the simulation were met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 3618 (Execution Processes for the Outer Planet UV Emissions: Future - Cycle C) 3.2 Complete PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.3 Complete FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: cycle 3 High) 3.4 Complete FOS RD 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 3.5 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.6 Complete Three Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.7 Complete FOC/96 4113 (Auroral Imaging of Jupiter with the FOC Cycle 2/ February-March 1993) 3.8 Complete Two Set of FOC/96 4821 (MAGELLANIC Cloud Planetary Nebulae: Cycle 3 Med.) 3.9 Complete Two Set of WFC 4849 (GTO High Latitude Parallel Program: Cycle 3 High) 3.10 Complete GHRS/2 4876 (Weak Interstellar Absorption Toward Zeta Orp - Augmentation Cycle 3 High) 3.11 Complete Uplink Updated Ephemeris Tables, Per MOSES 2871 3.12 Complete GHRS/2 4272 (Supersonic Chromoshperic Winds - Cycle 3 Med) 3.13 Complete Two Sets of FOS/BL 3545 (The Evolution of the UV Spectra in Early Type Galaxies Out to Z = 0.7) 3.14 Complete FOC/96 4969 (Search for QSOs Suitable for Subsequent Observations of He II) 3.15 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies with the FOC/96) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-